# Impact of cavities on the detection of quadratically coupled ultra-light dark matter

### v1

@@ IV.3 Discussion of Approximations
$m>6.58\times 10^{-16}\left(\frac{\mbox{s}}{\tau}\right)\leavevmode\nobreak\ \mbox{eV}\;,$ (41)
allow for a robust prediction of the $5^{\rm th}$ force. For experiments, such as MICROSCOPE and Eöt-Wash, which operate over time periods $\tau\sim\mathcal{O}(\text{days})$, we require $m\gtrsim 10^{-20}\mbox{ eV}$. Clearly, shorter duration experiments will be more restrictive in this regard. For example, if $\tau\sim\mathcal{O}(1\leavevmode\nobreak\ \text{s})$, our calculations can only be trusted empirically for $m\gtrsim 10^{-15}\leavevmode\nobreak\ \mbox{eV}$.
@@
$\left|\frac{B_{0}/L}{a_{0}}\right|\ll 1\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \left|\frac{B_{1}/L^{2}}{a_{1}L}\right|\ll 1\;,$ (42)
The linearisation of $\varphi_{\oplus}$ outside of the Earth, about $r_{c}$, holds for $r\ll r_{c}$, motivating $L=r_{c}$. However, the field profile due to the Earth, may vary significantly at the Earth’s surface, due to the rapid change in density, and therefore the effective mass of the scalar. Therefore, our linearisation of the field profile is only valid away from the surface. As such, we choose $L=\min(r_{c},\leavevmode\nobreak\ h)=h$ where as before, $h$ is the height of the centre of the cavity above the Earth’s surface.
?figurename? 7: A linearised background field profile is valid as long as the conditions in Eq. (42) are satisfied. We plot the relevant values on these conditions for cavity $1$ at $h=100\mbox{ m},\leavevmode\nobreak\ 10\mbox{ km}\leavevmode\nobreak\ \mbox{\leavevmode\nobreak\ and}\leavevmode\nobreak\ 1000\mbox{ km}$ (blue, red and green solid lines respectively) above the Earth. The left panels show the quantity in Eq. (42) corresponding to the condition on the degree 00 part of the field profile, and the panels on the right to degree $1$ part - displayed as a function of $\alpha$. When these functions are much smaller than one the linearised approximation is valid. The top row shows values for $\alpha>0$, and the bottom for $\alpha<0$.
[…]
@@
$\vec{F}_{5,G}=-\int d^{3}x\leavevmode\nobreak\ \nabla\Phi\rho(x)\alpha\frac{\phi^{2}}{2}\;.$ (43)
@@
$\phi(\vec{x},t)\approx\phi_{\rm st}(r,t)\left(1+\frac{m(kR-\tanh(kR))}{k}\right)\sqrt{\frac{2}{\pi mt}}$ (45)
where the ‘static’ profile $\phi_{\rm st}$ is the response to a static object embedded in the time-dependent cosmological background. It can be shown that the non-trivial $m$ and $\alpha$ dependence in Eq. (45) mean that, in the parameter space of interest, the field is always able to respond on the timescales of the motion of matter sources relevant to the experiment  [72]. Although these results were derived in the context of an instantaneously appearing solid sphere, we expect the same scaling of the response of the field with time to apply for the matter distributions we consider in this work. By a change of frame, we also expect Eq. (45) to indicate the relaxation time for a a static source in a time-dependent dark matter background, for example in the presence of a constant dark matter wind. We note that, in Refs. [52, 53], it was shown that dark matter plane waves impinging on an object from infinity can remove the divergences seen in the $\alpha<0$ case.
@@ V Conclusions
Quadratically coupled, ultra-light scalar fields are a well-motivated dark matter candidate. Their phenomenology has distinct differences to that of linearly coupled ultra-light scalars, largely driven by a density dependent effective mass for the scalar. In this work we have shown that hollow enclosures, or cavities, can suppress the amplitude of scalar oscillations within them, and that this suppression is particularly effective at strong couplings. We find this effect for both spherical and cylindrical cavities, and for positive and negative choices of coupling. In particular, the suppression is exponential in the case of positive couplings and follows a power law for negative couplings. In the case of negative couplings we also find that the field may diverge for discrete choices of coupling constant, depending on the extent of the matter distribution. Such divergences have been observed previously for solid sources, and require knowledge of the UV completion of the theory to resolve. The suppression of both amplitudes and gradients of the scalar field makes it challenging to detect the scalar field, and means that existing constraints at strong coupling should be reevaluated. Precise reevaluation likely requires detailed numerical modelling of experimental apparatus, and is beyond the scope of this work, however we illustrate the way in which existing constraints could be modified in Figure 5.
@@ Acknowledgements.
We would like to thank Benjamin Elder, Yeray Garcia del Castillo, and Joerg Jaeckel for interesting discussions and collaboration on related topics. CB and ET are supported by STFC Consolidated Grant [Grant No. ST/T000732/1]. AM is supported by EPSRC Quantum Technologies Doctoral Training Partnership (DTP) 2024-25. MR is supported by funding from the U.S. National Science Foundation under Awards PHY-1912380 and PHY-1912514. GR is supported by U.S. Department of Energy Grant No. DE-SC0011665. For open access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.
Data Availability Statement
In case of legitimate interest, the authors will provide the Python codes used for producing the plots in this work on request.
?refname?
@@
?appendixname? A Spherical Cavity Analytic Field Coefficients
@@
$\varphi_{c}(r)=\begin{cases}\sum_{l=0}^{\infty}\left(A_{l}^{in}r^{l}+B_{l}^{in}r^{-l-1}\right)P_{l}^{0}(\cos\theta)&r<R_{1}\\ \sum_{l=0}^{\infty}\left(\begin{split}&A_{l}^{walls}[(-i)^{l}j_{l}(-ikr)]\\ +&\leavevmode\nobreak\ B_{l}^{walls}[(-i)^{-l-1}y_{l}(-ikr)]\end{split}\right)P_{l}^{0}(\cos\theta)&R_{1}<r<R_{2}\\ \sum_{l=0}^{\infty}\left(A_{l}^{out}r^{l}+B_{l}^{out}r^{-l-1}\right)P_{l}^{0}(\cos\theta)&r>R_{2}\end{cases}$ (46)
where $A_{l}$ and $B_{l}$ are constants of integration, $k=\sqrt{\rho_{c}\alpha}$ and $i$ is the imaginary unit. $P^{0}_{l}(\cos\theta)$ are 0th order Legendre polynomials of $l^{th}$ degree, and $j_{l}(x),\leavevmode\nobreak\ y_{l}(x)$ are Spherical Bessel functions of the first and second kind respectively. The choice to include the factors $(-i)^{p}$ in front of the Bessel functions ensures that the coefficients $A_{l}^{walls}$, $B_{l}^{walls}$ are real for $k\in\mathbb{R}$. A valid solution requires continuity and differentiability at $R_{1}$ and $R_{2}$, regularity at the origin, and that the field approaches the background solution far from the cavity. Amalgamating these conditions, we ascribe the following expressions to the constants, noting that all but the degree 00 and degree $1$ terms vanish:
@@
?appendixname? B Solution for an infinitely long cylindrical cavity
@@
$\partial_{r}^{2}\varphi(r)+\frac{1}{r}\partial_{r}\varphi(r)=\alpha\rho(r)\varphi(r)\enspace,$ (57)
@@
$\rho(r)=\rho_{c}\,\Theta(r-R_{1})\Theta(R_{2}-r)\enspace.$ (58)
@@
?appendixname? C Cavity $5^{\rm{th}}$ Forces
@@
$\vec{F_{5}}=\int d^{3}x\leavevmode\nobreak\ \nabla\left(\rho(x)\alpha\right)\frac{\phi^{2}}{2}=-\int d^{3}x\leavevmode\nobreak\ \rho(x)\alpha\frac{\nabla\phi^{2}}{2}$ (66)
@@
$\vec{F}_{5}=-\frac{1}{2}\alpha\rho_{c}\left[\int_{R_{2}}d^{3}x\leavevmode\nobreak\ \nabla\phi^{2}-\int_{R_{1}}d^{3}x\leavevmode\nobreak\ \nabla\phi^{2}\right]$ (67)
@@
$\vec{F}_{5}=-\frac{1}{2}\alpha\rho_{c}\left[\int_{R_{2}}d^{2}\Omega\leavevmode\nobreak\ R_{2}^{2}\leavevmode\nobreak\ \hat{r}\leavevmode\nobreak\ \phi^{2}-\int_{R_{1}}d^{2}\Omega\leavevmode\nobreak\ R_{1}^{2}\leavevmode\nobreak\ \hat{r}\leavevmode\nobreak\ \phi^{2}\right]$ (68)
@@
$\begin{split}\vec{F}_{5}=-\pi\alpha\rho_{c}&\left[\int_{0}^{\pi}d\theta\sin\theta\leavevmode\nobreak\ R_{2}^{2}\leavevmode\nobreak\ (\hat{z}\cos\theta)\leavevmode\nobreak\ \left(a_{0}+\frac{B^{out}_{0}}{R_{2}}+\left(a_{1}R_{2}+\frac{B^{out}_{1}}{R_{2}^{2}}\right)\cos\theta\right)^{2}\right.\\ &\left.-\int_{0}^{\pi}d\theta\sin\theta\leavevmode\nobreak\ R_{1}^{2}\leavevmode\nobreak\ (\hat{z}\cos\theta)\leavevmode\nobreak\ (A^{in}_{0}+A^{in}_{1}R_{1}\cos\theta)^{2}\right]\end{split}$ (69)

### v2

@@ IV.3 Discussion of Approximations
$m>6.58\times 10^{-16}\left(\frac{\mbox{s}}{\tau}\right)\penalty 10000\ \mbox{eV}\;,$ (41)
allow for a robust prediction of the $5^{\rm th}$ force. For experiments, such as MICROSCOPE and Eöt-Wash, which operate over time periods $\tau\sim\mathcal{O}(\text{days})$, we require $m\gtrsim 10^{-20}\mbox{ eV}$. Clearly, shorter duration experiments will be more restrictive in this regard. For example, if $\tau\sim\mathcal{O}(1\penalty 10000\ \text{s})$, our calculations can only be trusted empirically for $m\gtrsim 10^{-15}\penalty 10000\ \mbox{eV}$.
@@
$\left|\frac{B_{0}/L}{a_{0}}\right|\ll 1\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \left|\frac{B_{1}/L^{2}}{a_{1}L}\right|\ll 1\;,$ (42)
The linearisation of $\varphi_{\oplus}$ outside of the Earth, about $r_{c}$, holds for $r\ll r_{c}$, motivating $L=r_{c}$. However, the field profile due to the Earth, may vary significantly at the Earth’s surface, due to the rapid change in density, and therefore the effective mass of the scalar. Therefore, our linearisation of the field profile is only valid away from the surface. As such, we choose $L=\min(r_{c},\penalty 10000\ h)=h$ where as before, $h$ is the height of the centre of the cavity above the Earth’s surface.
Figure 8: A linearised background field profile is valid as long as the conditions in Eq. (42) are satisfied. We plot the relevant values on these conditions for cavity 11 at $h=100\mbox{ m},\penalty 10000\ 10\mbox{ km}\penalty 10000\ \mbox{\penalty 10000\ and}\penalty 10000\ 1000\mbox{ km}$ (blue, red and green solid lines respectively) above the Earth. The left panels show the quantity in Eq. (42) corresponding to the condition on the degree 0 part of the field profile, and the panels on the right to degree 11 part - displayed as a function of $\alpha$. When these functions are much smaller than one the linearised approximation is valid. The top row shows values for $\alpha>0$, and the bottom for $\alpha<0$.
Figure 8 shows that the conditions in Eq. (42) generally hold for cavity 11 at $h=100\penalty 10000\ \mbox{m}$, $10\mbox{ km}$, and $1000\mbox{ km}$. The approximations may break down due to cavity induced divergences when $\alpha<0$, though some level of enhancement is allowed. The conditions become harder to satisfy at smaller $h$, becoming completely invalid for $h\approx 0$. Therefore, for an experiment at the surface of the Earth, the use of numerical methods will likely be required to fully model the scalar field profile. However, we expect that intra-cavity suppression and enhancement of the scalar field, discussed in Section III.2, should should still be seen.
@@
$\vec{F}_{5,G}=-\int d^{3}x\penalty 10000\ \nabla\Phi\rho(x)\alpha\frac{\phi^{2}}{2}\;.$ (43)
@@
$\phi(\vec{x},t)\approx\phi_{\rm st}(r,t)\left(1+\frac{m(kR-\tanh(kR))}{k}\sqrt{\frac{2}{\pi mt}}\right)\;,$ (45)
where the ‘static’ profile $\phi_{\rm st}$ is the response to a static object embedded in the time-dependent cosmological background. It can be shown that the non-trivial $m$ and $\alpha$ dependence in Eq. (45) mean that, in the parameter space of interest, the field is always able to respond on the timescales of the motion of matter sources relevant to the experiment [26]. Although these results were derived in the context of an instantaneously appearing solid sphere, we expect the same scaling of the response of the field with time to apply for the matter distributions we consider in this work. By a change of frame, we also expect Eq. (45) to indicate the relaxation time for a static source in a time-dependent dark matter background, for example in the presence of a constant dark matter wind. We note that, in Refs. [13, 35], it was shown that dark matter plane waves impinging on an object from infinity can remove the divergences seen in the $\alpha<0$ case.
@@ V Conclusions
Quadratically coupled, ultra-light scalar fields are a well-motivated dark matter candidate. Their phenomenology has distinct differences to that of linearly coupled ultra-light scalars, largely driven by a density dependent effective mass for the scalar. In this work we have shown that hollow enclosures, or cavities, can suppress the amplitude of scalar oscillations within them, and that this suppression is particularly effective at strong couplings. We find this effect for both spherical and cylindrical cavities, and for positive and negative choices of coupling. In particular, the suppression is exponential in the case of positive couplings and follows a power law trend for negative couplings. In the case of negative couplings, we also find that the field amplitude may diverge for discrete choices of coupling constant, which depend on the extent of the matter distribution. Such divergences have been observed previously for solid sources. Enhancement of the scalar field is predicted when the coupling constant is in the neighbourhood of these discrete values. However, where the expected enhancement is too great, knowledge of the UV completion of the scalar theory is required to resolve the true behaviour of the field. The general trends of suppression in both amplitude and gradient of the scalar profile makes it challenging to detect the scalar field, and means that existing constraints at strong coupling should be reevaluated. Precise reevaluation likely requires detailed numerical modelling of experimental apparatus, and is beyond the scope of this work, however we illustrate the way in which existing constraints could be modified in Figure 6.
@@ Acknowledgements.
We would like to thank Benjamin Elder, Yeray Garcia del Castillo, and Joerg Jaeckel for interesting discussions and collaboration on related topics. CB and ET are supported by STFC Consolidated Grant [Grant No. ST/T000732/1]. AM is supported by EPSRC Quantum Technologies Doctoral Training Partnership (DTP) 2024-25. MR is supported by funding from the U.S. National Science Foundation under Awards PHY-1912380 and PHY-1912514. GR is supported by U.S. Department of Energy Grant No. DE-SC0011665. For open access purposes, the author has applied a $C$ BY public copyright license to any author accepted manuscript version arising from this submission.
@@
Appendix A Spherical Cavity Analytic Field Coefficients
@@
$\varphi_{c}(r)=\begin{cases}\sum_{l=0}^{\infty}\left(A_{l}^{in}r^{l}+B_{l}^{in}r^{-l-1}\right)P_{l}^{0}(\cos\theta)&r<R_{1}\\ \sum_{l=0}^{\infty}\left(\begin{split}&A_{l}^{walls}[(-i)^{l}j_{l}(-ikr)]\\ +&\penalty 10000\ B_{l}^{walls}[(-i)^{-l-1}y_{l}(-ikr)]\end{split}\right)P_{l}^{0}(\cos\theta)&R_{1}<r<R_{2}\\ \sum_{l=0}^{\infty}\left(A_{l}^{out}r^{l}+B_{l}^{out}r^{-l-1}\right)P_{l}^{0}(\cos\theta)&r>R_{2}\end{cases}$ (46)
where $A_{l}$ and $B_{l}$ are constants of integration, $k=\sqrt{\rho_{c}\alpha}$ and $i$ is the imaginary unit. $P^{0}_{l}(\cos\theta)$ are 0th order Legendre polynomials of $l^{th}$ degree, and $j_{l}(x),\penalty 10000\ y_{l}(x)$ are Spherical Bessel functions of the first and second kind respectively. The choice to include the factors $(-i)^{p}$ in front of the Bessel functions ensures that the coefficients $A_{l}^{walls}$, $B_{l}^{walls}$ are real for $k\in\mathbb{R}$. A valid solution requires continuity and differentiability at $R_{1}$ and $R_{2}$, regularity at the origin, and that the field approaches the background solution far from the cavity. Amalgamating these conditions, we ascribe the following expressions to the constants, noting that all but the degree 0 and degree 11 terms vanish:
@@
Appendix B Solution for an infinitely long cylindrical cavity
@@
$\partial_{r}^{2}\varphi(r)+\frac{1}{r}\partial_{r}\varphi(r)=\alpha\rho(r)\varphi(r)\kern 5.0pt,$ (57)
@@
$\rho(r)=\rho_{c}\,\Theta(r-R_{1})\Theta(R_{2}-r)\kern 5.0pt.$ (58)
@@
Appendix C Cavity $5^{\rm{th}}$ Forces
@@
$\vec{F_{5}}=\int d^{3}x\penalty 10000\ \nabla\left(\rho(x)\alpha\right)\frac{\phi^{2}}{2}=-\int d^{3}x\penalty 10000\ \rho(x)\alpha\frac{\nabla\phi^{2}}{2}$ (66)
@@
$\vec{F}_{5}=-\frac{1}{2}\alpha\rho_{c}\left[\int_{R_{2}}d^{3}x\penalty 10000\ \nabla\phi^{2}-\int_{R_{1}}d^{3}x\penalty 10000\ \nabla\phi^{2}\right]$ (67)
@@
$\vec{F}_{5}=-\frac{1}{2}\alpha\rho_{c}\left[\int_{R_{2}}d^{2}\Omega\penalty 10000\ R_{2}^{2}\penalty 10000\ \hat{r}\penalty 10000\ \phi^{2}-\int_{R_{1}}d^{2}\Omega\penalty 10000\ R_{1}^{2}\penalty 10000\ \hat{r}\penalty 10000\ \phi^{2}\right]$ (68)
@@
$\begin{split}\vec{F}_{5}=-\pi\alpha\rho_{c}&\left[\int_{0}^{\pi}d\theta\sin\theta\penalty 10000\ R_{2}^{2}\penalty 10000\ (\hat{z}\cos\theta)\penalty 10000\ \left(a_{0}+\frac{B^{out}_{0}}{R_{2}}+\left(a_{1}R_{2}+\frac{B^{out}_{1}}{R_{2}^{2}}\right)\cos\theta\right)^{2}\right.\\ &\left.-\int_{0}^{\pi}d\theta\sin\theta\penalty 10000\ R_{1}^{2}\penalty 10000\ (\hat{z}\cos\theta)\penalty 10000\ (A^{in}_{0}+A^{in}_{1}R_{1}\cos\theta)^{2}\right]\end{split}$ (69)